# Representations of a quantum-deformed Lorentz algebra, Clebsch-Gordan map, and Fenchel-Nielsen representation of quantum complex flat connections at level-$\bm{k}$

### v1

@@ 5.1 Periodic states
Inserting (5.28) in (5.27), we check that the integrand of $\int\mathrm{d}\mu\mathrm{d}s$ is a Schwarz function on $\mathbb{R}^{2}$  [^9] The integrand suppresses exponentially fast as $\mu,\mathrm{Re}(s)\to\infty$, as can be shown by using ${\rm D}_{b}(\frac{c_{b}}{\sqrt{N}}+x,n)\sim\exp\left[i\pi\left(x+\frac{c_{b}}{\sqrt{N}}\right)^{2}+O(1)\right]$ as $\mathrm{Re}(x)\to\infty$ and ${\rm D}_{b}(\frac{c_{b}}{\sqrt{N}}+x,n)\sim O(1)$ as $\mathrm{Re}(x)\to-\infty$ [12]. , so we can interchange the order of integration. Then we compute the integral and sum of $\mu=x\sqrt{N},m\in\mathbb{Z}/N\mathbb{Z}$ by applying (A.15) 101010 We replace the variables in (A.15) by $t\to x,\ \alpha\to-\lambda^{\prime}-\omega^{\prime\prime}+2i\epsilon,\ \beta\to-\lambda+\omega^{\prime\prime}-2i\epsilon,\ s\to-s+2i\epsilon-2\omega^{\prime\prime},\ a\to-m_{r}^{\prime}/2,\ p\to-m_{r}/2,\ c\to-c$ (the left-hand sides are variables in (A.15), while the right-hand sides are variables in (5.28)). We have $\mathrm{Im}\left(\alpha+\frac{c_{b}}{\sqrt{N}}\right)\to\epsilon>0,\ \mathrm{Im}\left(\frac{c_{b}}{\sqrt{N}}-\beta\right)\to\epsilon>0,\ \mathrm{Im}(\alpha-\beta)\to 2\epsilon-\frac{b+b^{-1}}{\sqrt{N}}<0$ for sufficiently small $\epsilon>0$, and $\mathrm{Im}(s)\to 2\epsilon-\frac{b+b^{-1}}{\sqrt{N}}-\mathrm{Im}(s)$. Therefore for sufficiently small $\epsilon>0$, $\mathrm{Im}(s)<0$ and small $|\mathrm{Im}(s)|$ in (5.28) implies $\mathrm{Im}(\alpha-\beta)<\mathrm{Im}(s)<0$ in (A.15), ensuring the validity of (A.15). then using the inverse formula and recusion relation of $\gamma$. We obtain that as $\epsilon\to 0$
@@
$\sum_{m\in\mathbb{Z}/N\mathbb{Z}}\int\mathrm{d}\mu\,\psi_{r}(\mu,m)^{*}\psi_{r^{\prime}}(\mu,m)$ $=$ $\frac{1}{2\sqrt{N}}\rho(\mu_{r},m_{r})\sum_{c\in\mathbb{Z}/N\mathbb{Z}}\int\mathrm{d}s\,e^{-\frac{2\pi i}{N}c\left(\frac{m_{r^{\prime}}}{2}+\frac{m_{r}}{2}\right)+2i\pi s\left(\lambda^{\prime}+\lambda\right)}$ (5.29)
$=$ $\frac{1}{2}\rho(\mu_{r},m_{r})\delta(\mu_{r}+\mu_{r^{\prime}})\delta_{m_{r}+m_{r^{\prime}},0}\ .$
@@
This equation holds in the sense of tempered distribution [^11] The result of $s$-integral is understood as a tempered distribution, which makes sense only when interring another integral, say, over $\mu_{r}=2\sqrt{N}\lambda_{r}$. For finite $\epsilon$, the integrand of $\int\mathrm{d}s$ is $e^{-\frac{2\pi i}{N}c\left(\frac{m_{r^{\prime}}}{2}+\frac{m_{r}}{2}\right)+2i\pi s\left(\lambda^{\prime}+\lambda\right)}$ multiply a Schwarz function in $s$. Consider the following integral: $\int\mathrm{d}x\mathrm{d}k\ e^{2\pi ik(x-x^{\prime})}f_{\epsilon}(k)F(x)$ with both $f_{\epsilon},F$ Schwarz functions and $\lim_{\epsilon\to 0}f_{\epsilon}(k)=1$. This integral equals $\int\mathrm{d}k\,e^{-2\pi ikx^{\prime}}f_{\epsilon}(k)\widetilde{F}(k)=f(x^{\prime})$ as $\epsilon\to 0$, where the limit and integral can be interchanged by the dominant convergence. The computation in (5.29) is the same.. Finally, using the reflection symmetry of $\psi_{r}$
$\sum_{m\in\mathbb{Z}/k\mathbb{Z}}\int\mathrm{d}\mu\,\psi_{r}(\mu,m)^{*}\psi_{r^{\prime}}(\mu,m)$ $=$ $\frac{1}{2}\sum_{m\in\mathbb{Z}/k\mathbb{Z}}\int\mathrm{d}\mu\,\psi_{r}(\mu,m)^{*}\left[\psi_{r^{\prime}}(\mu,m)+\psi_{r^{\prime}{}^{-1}}(\mu,m)\right]$ (5.30)
$=$ $\rho(\mu_{r},m_{r})\,\delta_{r,r^{\prime}}$
@@
The eigenstates satisfy the following resolution of identity on $\mathcal{H}_{+}$
$2\sum_{m_{r}/2\in\mathbb{Z}/N\mathbb{Z}}\int_{0}^{\infty}\mathrm{d}\mu_{r}\,\rho(\mu_{r},m_{r})^{-1}\,\psi_{r}(\mu^{\prime},m^{\prime})\,\psi_{r}(\mu,m)^{*}=\delta(\mu,\mu^{\prime})\,\delta_{m,m^{\prime}}.$ (5.31)
@@
$\sigma(\mu_{r},m_{r})=\frac{e^{-4i\pi\lambda\omega^{\prime\prime}}e^{-\frac{i\pi m_{r}}{N}+4i\pi m_{r}\omega^{\prime}{}^{2}-4i\pi m_{r}\omega^{\prime}{}\omega^{\prime\prime}}-e^{-4i\pi\lambda(\omega^{\prime\prime}-2\omega^{\prime}{})}e^{\frac{i\pi m_{r}}{N}+4i\pi m_{r}\omega^{\prime}{}^{2}-4i\pi m_{r}\omega^{\prime}{}\omega^{\prime\prime}}}{4N^{2}},$ (5.32)
@@
$\sum_{m_{r}/2\in\mathbb{Z}/N\mathbb{Z}}\int_{-\infty}^{\infty}\mathrm{d}\mu_{r}\,\sigma(\mu_{r},m_{r})\,\psi_{r}(\mu^{\prime},m^{\prime})\,\psi_{r}(\mu,m)^{*}$ $=$ $\sum_{m_{r}/2\in\mathbb{Z}/N\mathbb{Z}}\int_{0}^{\infty}\mathrm{d}\mu_{r}\,\rho(\mu_{r},m_{r})^{-1}\,\psi_{r}(\mu^{\prime},m^{\prime})\,\psi_{r}(\mu,m)^{*}.$
@@
$I$ $=$ $\sum_{m_{r}/2\in\mathbb{Z}/N\mathbb{Z}}\int_{-\infty}^{\infty}\mathrm{d}\mu_{r}\,\sigma(\mu_{r},m_{r})\,\psi_{r}(\mu^{\prime},m^{\prime})\,\psi_{r}(\mu,m)^{*}$ (5.33)
$=$ $2\sqrt{N}\sum_{m_{r}/2\in\mathbb{Z}/N\mathbb{Z}}\int_{-\infty}^{\infty}\mathrm{d}\lambda\,\sigma\,e^{-\frac{i\pi\left(m^{2}-{m^{\prime}}^{2}\right)}{N}+i\pi(m-{m^{\prime}})+i\pi\left(x^{2}-y^{2}\right)+2i\pi(x+y)\omega^{\prime\prime}}$
$\frac{\gamma\left(-\lambda-\omega^{\prime\prime}+y+i\epsilon,m^{\prime}-\frac{m_{r}}{2}\right)\gamma\left(\lambda-\omega^{\prime\prime}+y+i\epsilon,\frac{m_{r}}{2}+m^{\prime}\right)}{\gamma\left(-\lambda+x+\omega^{\prime\prime}-i\epsilon,m-\frac{m_{r}}{2}\right)\gamma\left(\lambda+x+\omega^{\prime\prime}-i\epsilon,\frac{m_{r}}{2}+m\right)},$
@@
$-t=\lambda,\qquad\alpha=-\omega^{\prime\prime}+y+i\epsilon,\qquad\beta=x+\omega^{\prime\prime}-i\epsilon,\qquad a=m^{\prime},\qquad p=m,\qquad-d=\frac{m_{r}}{2}.$ (5.34)
Then we apply (A.17) for $\mathrm{Im}(\alpha+\frac{c_{b}}{\sqrt{N}})=\mathrm{Im}(-\beta+\frac{c_{b}}{\sqrt{N}})=\epsilon>0$ and $\mathrm{Im}(\alpha-\beta)=2\epsilon-\frac{2\mathrm{Im}(c_{b})}{\sqrt{N}}<\mathrm{Im}(s)<0$. The result is
@@
$I_{1}$ $=$ $\frac{\zeta_{0}}{2N^{2}}\sum_{d,c\in\mathbb{Z}/N\mathbb{Z}}\int\mathrm{d}t\mathrm{d}s\,e^{-i\pi\left(\alpha^{2}-\beta^{2}\right)+\frac{i\pi\left(a^{2}+2ac-p^{2}\right)}{N}-i\pi(a-p)-\frac{2i\pi}{N}dc}$ (5.36)
@@
$I_{2}$ $=$ $-\frac{\zeta_{0}}{2N^{2}}\sum_{d,c\in\mathbb{Z}/N\mathbb{Z}}\int\mathrm{d}t\mathrm{d}s\,e^{-i\pi\left(\alpha^{2}-\beta^{2}\right)+\frac{i\pi\left(a^{2}+2ac-p^{2}\right)}{N}-i\pi(a-p)-\frac{2i\pi}{N}d\left(c+2\right)}$ (5.37)
@@
$I_{1}$ $=$ $\frac{\zeta_{0}^{2}}{2N^{3/2}}\sum_{c\in\mathbb{Z}/N\mathbb{Z}}\int\mathrm{d}s\,e^{\frac{4i\pi cm^{\prime}{}}{N}-\frac{i\pi\left(m^{2}-m^{\prime}{}^{2}\right)}{N}+i\pi(m-m^{\prime}{})+4i\pi sy+i\pi\left(x^{2}-y^{2}\right)+2i\pi(x-y)\omega^{\prime\prime}-2\pi\delta y}$ (5.38)
@@
$I_{2}$ $=$ $-\frac{\zeta_{0}^{2}}{2N^{3/2}}\sum_{c\in\mathbb{Z}/N\mathbb{Z}}\int\mathrm{d}s\,e^{\frac{2i\pi(2cm^{\prime}{}+2m^{\prime}{})}{N}-\frac{i\pi\left(m^{2}-m^{\prime}{}^{2}\right)}{N}+i\pi(m-m^{\prime}{})+4i\pi sy+i\pi\left(x^{2}-y^{2}\right)+2i\pi(x+3y)\omega^{\prime}+2i\pi\omega(x-y)}$ (5.39)
@@
The poles and zeros of $D_{b}(x,n)=\gamma(-x,n)$ ($\mathrm{Im}(b)>0$) is respectively at $x=\frac{1}{\sqrt{N}}({c_{b}}+ib^{-1}l+ibm)$ for $n=m-l+N\mathbb{Z}$ and $x=-\frac{1}{\sqrt{N}}({c_{b}}+ib^{-1}l+ibm)$ for $n=l-m+N\mathbb{Z}$, where $l,m\geq 0$ for both [11]. The integrand of $I_{1}$ has the following poles ($l,m\geq 0$ for below):
@@
and ${\rm D}_{\mathrm{b}}\left(-\frac{c_{b}}{\sqrt{N}}+s,0\right)\sim-\frac{2\pi b}{\sqrt{N}}s(\mathbf{q}^{2},\mathbf{q}^{2})_{\infty}/(\tilde{\mathbf{q}}^{-2},\tilde{\mathbf{q}}^{-2})_{\infty}$ as $s\to 0$. The residue of $I_{1}$ at $s_{\rm pole}=0$ vanishes as $\delta\to 0$ unless $m=m^{\prime}$ As a result, for $\delta,\epsilon\to 0$ with $\delta>2\epsilon$, we obtain
$I=\delta_{m,m^{\prime}}\frac{\delta}{4\pi\sqrt{N}(x-y-2i\epsilon)(i(\delta-2\epsilon)+x-y)}\to\frac{1}{2}\delta(\mu-\mu^{\prime})\delta_{m,m^{\prime}}.$ (5.45)
The result indicates that for any $f_{+}\in\mathcal{H}_{+}$, we can find the spectral representation $f(\mu_{r},m_{r})$ with $\mu_{r}\in\mathbb{R}_{+}$, $m_{r}/2\in\mathbb{Z}/N\mathbb{Z}$,
$\mid f_{+}\rangle$ $=$ $\sum_{m_{r}/2\in\mathbb{Z}/N\mathbb{Z}}\int_{0}^{\infty}\mathrm{d}\mu_{r}\,\rho(\mu_{r},m_{r})^{-1}f_{+}(\mu_{r},m_{r})\mid\psi_{r}\rangle,$
where $\quad f_{+}(\mu_{r},m_{r})=\langle\psi_{r}\mid f_{+}\rangle=2\langle\psi_{r}\mid f_{+}\rangle_{N}$ (5.46)
@@
$\langle f_{+}\mid f_{+}^{\prime}\rangle=\sum_{m_{r}/2\in\mathbb{Z}/N\mathbb{Z}}\int_{0}^{\infty}\mathrm{d}\mu_{r}\,\rho(\mu_{r},m_{r})^{-1}f_{+}(\mu_{r},m_{r})^{*}f^{\prime}_{+}(\mu_{r},m_{r}).$ (5.47)
@@
$\bm{L}f_{+}(\mu_{r},m_{r})=\ell(r)f_{+}(\mu_{r},m_{r}),\qquad\tilde{\bm{L}}f_{+}(\mu_{r},m_{r})=\ell(r)^{*}f_{+}(\mu_{r},m_{r}).$ (5.48)
The spectral decomposition can be conveniently represented as the direct integral decomposition
$\mathcal{H}_{+}\simeq\sum_{m_{r}/2\in\mathbb{Z}/N\mathbb{Z}}\int_{0}^{\infty}\mathrm{d}\mu_{r}\,\rho(\mu_{r},m_{r})^{-1}\,\mathcal{H}_{\mu_{r},m_{r}}.$ (5.49)

### v2

@@ 5 Spectral decomposition
Inserting (5.12) in (5.11), we check that the integrand of $\int\mathrm{d}\mu\mathrm{d}s$ is a Schwarz function on $\mathbb{R}^{2}$ for $\epsilon>0$  [^6] The integrand suppresses exponentially fast as $\mu,\mathrm{Re}(s)\to\infty$, as can be shown by using ${\rm D}_{b}(\frac{c_{b}}{\sqrt{N}}+x,n)\sim\exp\left[i\pi\left(x+\frac{c_{b}}{\sqrt{N}}\right)^{2}+O(1)\right]$ as $\mathrm{Re}(x)\to\infty$ and ${\rm D}_{b}(\frac{c_{b}}{\sqrt{N}}+x,n)\sim O(1)$ as $\mathrm{Re}(x)\to-\infty$ [12]. , so we can interchange the order of integration. Then we compute the integral and sum of $\mu=x\sqrt{N},m\in\mathbb{Z}/N\mathbb{Z}$ by applying (A.16) 777 We replace the variables in (A.16) by $t\to x,\ \alpha\to-\lambda^{\prime}-\omega^{\prime\prime}+2i\epsilon,\ \beta\to-\lambda+\omega^{\prime\prime}-2i\epsilon,\ s\to-s+2i\epsilon-2\omega^{\prime\prime},\ a\to-m_{r}^{\prime},\ p\to-m_{r},\ c\to-c$ (the left-hand sides are variables in (A.16), while the right-hand sides are variables in (5.12)). We have $\mathrm{Im}\left(\alpha+\frac{c_{b}}{\sqrt{N}}\right)\to\epsilon>0,\ \mathrm{Im}\left(\frac{c_{b}}{\sqrt{N}}-\beta\right)\to\epsilon>0,\ \mathrm{Im}(\alpha-\beta)\to 2\epsilon-\frac{b+b^{-1}}{\sqrt{N}}<0$ for sufficiently small $\epsilon>0$, and $\mathrm{Im}(s)\to 2\epsilon-\frac{b+b^{-1}}{\sqrt{N}}-\mathrm{Im}(s)$. Therefore for sufficiently small $\epsilon>0$, $\mathrm{Im}(s)<0$ and small $|\mathrm{Im}(s)|$ in (5.12) implies $\mathrm{Im}(\alpha-\beta)<\mathrm{Im}(s)<0$ in (A.16), ensuring the validity of (A.16). then using the inverse formula and recusion relation of $\gamma$. We obtain that as $\epsilon\to 0$, the result has the singularities at $(\mu_{r^{\prime}},m_{r^{\prime}})=\pm(\mu_{r},m_{r})$. Away from $(\mu_{r^{\prime}},m_{r^{\prime}})=(\mu_{r},m_{r})$,
@@
$\sum_{m\in\mathbb{Z}/N\mathbb{Z}}\int\mathrm{d}\mu\,\psi_{r}(\mu,m)^{*}\psi_{r^{\prime}}(\mu,m)$ $=$ $\frac{1}{4\sqrt{N}}\rho(\mu_{r},m_{r})\sum_{c\in\mathbb{Z}/N\mathbb{Z}}\int\mathrm{d}s\,e^{-\frac{2\pi i}{N}c\left({m_{r^{\prime}}}+{m_{r}}\right)+2i\pi s\left(\lambda^{\prime}+\lambda\right)}$ (5.13)
$=$ $\frac{1}{4}\rho(\mu_{r},m_{r})\delta(\mu_{r}+\mu_{r^{\prime}})\delta_{m_{r}+m_{r^{\prime}},0}\ .$
@@
This equation holds in the sense of tempered distribution [^8] The result of $s$-integral is understood as a tempered distribution, which makes sense only when interring another integral, say, over $\mu_{r}=\sqrt{N}\lambda_{r}$. For finite $\epsilon$, the integrand of $\int\mathrm{d}s$ is $e^{-\frac{2\pi i}{N}c\left({m_{r^{\prime}}}+{m_{r}}\right)+2i\pi s\left(\lambda^{\prime}+\lambda\right)}$ multiplying a Schwarz function in $s$. Consider the following integral: $\int\mathrm{d}\lambda\mathrm{d}s\ e^{2\pi is(\lambda-\lambda^{\prime})}f_{\epsilon}(s)F(\lambda)$ with both $f_{\epsilon},F$ Schwarz functions and $\lim_{\epsilon\to 0}f_{\epsilon}(s)=1$. This integral equals $\int\mathrm{d}s\,e^{-2\pi is\lambda^{\prime}}f_{\epsilon}(s)\widetilde{F}(s)=F(\lambda^{\prime})$ as $\epsilon\to 0$, where the limit and integral can be interchanged by the dominant convergence. The computation in (5.13) is the same.. The behavior at the other singularity follows from the reflection symmetry of $\psi_{r}$
$\sum_{m\in\mathbb{Z}/N\mathbb{Z}}\int\mathrm{d}\mu\,\psi_{r}(\mu,m)^{*}\psi_{r^{\prime}}(\mu,m)=\frac{1}{4}\rho(\mu_{r},m_{r})\left[\delta(\mu_{r}-\mu_{r^{\prime}})\delta_{m_{r},m_{r^{\prime}}}+\delta(\mu_{r}+\mu_{r^{\prime}})\delta_{m_{r},-m_{r^{\prime}}}\right]$ (5.14)
@@
The eigenstates satisfy the following resolution of identity on $\mathcal{H}$
$4\sum_{m_{r}\in\mathbb{Z}/N\mathbb{Z}}\int_{0}^{\infty}\mathrm{d}\mu_{r}\,\rho(\mu_{r},m_{r})^{-1}\,\psi_{r}(\mu^{\prime},m^{\prime})\,\psi_{r}(\mu,m)^{*}=\delta(\mu,\mu^{\prime})\,\delta_{m,m^{\prime}}.$ (5.15)
@@
$\sigma(\mu_{r},m_{r})=\frac{e^{-4i\pi\lambda\omega^{\prime\prime}-\frac{2i\pi m_{r}}{N}+8i\pi m_{r}\omega^{\prime}{}^{2}-8i\pi m_{r}\omega^{\prime}\omega^{\prime\prime}}-e^{-4i\pi\lambda\left(\omega^{\prime\prime}-2\omega^{\prime}\right)+\frac{2i\pi m_{r}}{N}+8i\pi m_{r}\omega^{\prime}{}^{2}-8i\pi m_{r}\omega^{\prime}\omega^{\prime\prime}}}{4N^{2}},$
@@
$\sum_{m_{r}\in\mathbb{Z}/N\mathbb{Z}}\int_{-\infty}^{\infty}\mathrm{d}\mu_{r}\,\sigma(\mu_{r},m_{r})\,\psi_{r}(\mu^{\prime},m^{\prime})\,\psi_{r}(\mu,m)^{*}$ $=$ $\sum_{m_{r}\in\mathbb{Z}/N\mathbb{Z}}\int_{0}^{\infty}\mathrm{d}\mu_{r}\,\rho(\mu_{r},m_{r})^{-1}\,\psi_{r}(\mu^{\prime},m^{\prime})\,\psi_{r}(\mu,m)^{*}.$
@@
$I$ $=$ $\sum_{m_{r}\in\mathbb{Z}/N\mathbb{Z}}\int_{-\infty}^{\infty}\mathrm{d}\mu_{r}\,\sigma(\mu_{r},m_{r})\,\psi_{r}(\mu^{\prime},m^{\prime})\,\psi_{r}(\mu,m)^{*}$ (5.16)
$=$ $\sqrt{N}\sum_{m_{r}\in\mathbb{Z}/N\mathbb{Z}}\int_{-\infty}^{\infty}\mathrm{d}\lambda\,\sigma\,e^{-\frac{i\pi\left(m^{2}-{m^{\prime}}^{2}\right)}{N}+i\pi(m-{m^{\prime}})+i\pi\left(x^{2}-y^{2}\right)+2i\pi(x+y)(\omega^{\prime\prime}-i\epsilon)}$
$\frac{\gamma\left(-\lambda-\omega^{\prime\prime}+y+i\epsilon,m^{\prime}-{m_{r}}\right)\gamma\left(\lambda-\omega^{\prime\prime}+y+i\epsilon,{m_{r}}+m^{\prime}\right)}{\gamma\left(-\lambda+x+\omega^{\prime\prime}-i\epsilon,m-{m_{r}}\right)\gamma\left(\lambda+x+\omega^{\prime\prime}-i\epsilon,{m_{r}}+m\right)},$
@@
$-t=\lambda,\qquad\alpha=-\omega^{\prime\prime}+y+i\epsilon,\qquad\beta=x+\omega^{\prime\prime}-i\epsilon,\qquad a=m^{\prime},\qquad p=m,\qquad-d={m_{r}}.$
Then we apply (A.18) with $\mathrm{Im}(\alpha+\frac{c_{b}}{\sqrt{N}})=\mathrm{Im}(-\beta+\frac{c_{b}}{\sqrt{N}})=\epsilon>0$ and $\mathrm{Im}(\alpha-\beta)=2\epsilon-\frac{2\mathrm{Im}(c_{b})}{\sqrt{N}}<\mathrm{Im}(s)<0$. The result is
@@
$I_{1}$ $=$ $\frac{\zeta_{0}}{4N^{2}}\sum_{d,c\in\mathbb{Z}/N\mathbb{Z}}\int\mathrm{d}t\mathrm{d}s\,e^{-i\pi\left(\alpha^{2}-\beta^{2}\right)+\frac{i\pi\left(a^{2}+2ac-p^{2}\right)}{N}-i\pi(a-p)-\frac{2i\pi}{N}dc}$ (5.18)
@@
$I_{2}$ $=$ $-\frac{\zeta_{0}}{4N^{2}}\sum_{d,c\in\mathbb{Z}/N\mathbb{Z}}\int\mathrm{d}t\mathrm{d}s\,e^{-i\pi\left(\alpha^{2}-\beta^{2}\right)+\frac{i\pi\left(a^{2}+2ac-p^{2}\right)}{N}-i\pi(a-p)-\frac{2i\pi}{N}d\left(c+2\right)}$ (5.19)
@@
$I_{1}$ $=$ $\frac{\zeta_{0}^{2}}{4N^{3/2}}\sum_{c\in\mathbb{Z}/N\mathbb{Z}}\int\mathrm{d}s\,e^{\frac{4i\pi cm^{\prime}{}}{N}-\frac{i\pi\left(m^{2}-m^{\prime}{}^{2}\right)}{N}+i\pi(m-m^{\prime}{})+4i\pi sy+i\pi\left(x^{2}-y^{2}\right)+2i\pi(x-y)\omega^{\prime\prime}-2\pi\delta y}$ (5.20)
@@
$I_{2}$ $=$ $-\frac{\zeta_{0}^{2}}{4N^{3/2}}\sum_{c\in\mathbb{Z}/N\mathbb{Z}}\int\mathrm{d}s\,e^{\frac{2i\pi(2cm^{\prime}{}+2m^{\prime}{})}{N}-\frac{i\pi\left(m^{2}-m^{\prime}{}^{2}\right)}{N}+i\pi(m-m^{\prime}{})+4i\pi sy+i\pi\left(x^{2}-y^{2}\right)+2i\pi(x+3y)\omega^{\prime}+2i\pi\omega(x-y)}$ (5.21)
@@
The poles and zeros of $\mathrm{D}_{b}(x,n)=\gamma(-x,n)$ ($\mathrm{Im}(b)>0$) is respectively at $x=\frac{1}{\sqrt{N}}({c_{b}}+ib^{-1}l+ibm)$ for $n=m-l+N\mathbb{Z}$ and $x=-\frac{1}{\sqrt{N}}({c_{b}}+ib^{-1}l+ibm)$ for $n=l-m+N\mathbb{Z}$, where $l,m\geq 0$ for both [11]. The integrand of $I_{1}$ has the following poles ($l,m\geq 0$ for below):
@@
and ${\rm D}_{\mathrm{b}}\left(-\frac{c_{b}}{\sqrt{N}}+s,0\right)\sim-\frac{2\pi b}{\sqrt{N}}s(\mathbf{q}^{2},\mathbf{q}^{2})_{\infty}/(\tilde{\mathbf{q}}^{-2},\tilde{\mathbf{q}}^{-2})_{\infty}$ as $s\to 0$. The residue of $I_{1}$ at $s_{\rm pole}=0$ vanishes as $\delta\to 0$ unless $m=m^{\prime}$ and $x\to y$. As a result, for $\delta,\epsilon\to 0$ with $\delta>2\epsilon$, we obtain
$I=\delta_{m,m^{\prime}}\frac{\delta}{8\pi\sqrt{N}(x-y-2i\epsilon)(i(\delta-2\epsilon)+x-y)}\to\frac{1}{4}\delta(\mu-\mu^{\prime})\delta_{m,m^{\prime}}.$ (5.27)
For convenience, we denote by
$\varrho(\mu_{r},m_{r})=\frac{1}{4}\rho(\mu_{r},m_{r})=\frac{N^{2}}{4}\left[\sin\left(\frac{2\pi}{N}(ib\mu_{r}+m_{r})\right)\sin\left(\frac{2\pi}{N}(-ib^{-1}\mu_{r}+m_{r})\right)\right]^{-1}.$ (5.28)
The above result indicates that for any $f\in\mathcal{H}$, we can find the spectral representation $f(\mu_{r},m_{r})$ with $\mu_{r}\in\mathbb{R}_{+}$, $m_{r}\in\mathbb{Z}/N\mathbb{Z}$,
$\mid f\rangle$ $=$ $\sum_{m_{r}\in\mathbb{Z}/N\mathbb{Z}}\int_{0}^{\infty}\mathrm{d}\mu_{r}\,\varrho(\mu_{r},m_{r})^{-1}f(\mu_{r},m_{r})\mid\psi_{r}\rangle,$
where $\quad f(\mu_{r},m_{r})=\langle\psi_{r}\mid f\rangle.$ (5.29)
@@
$\langle f\mid f^{\prime}\rangle=\sum_{m_{r}\in\mathbb{Z}/N\mathbb{Z}}\int_{0}^{\infty}\mathrm{d}\mu_{r}\,\varrho(\mu_{r},m_{r})^{-1}f(\mu_{r},m_{r})^{*}f^{\prime}(\mu_{r},m_{r}).$ (5.30)
@@
where $\ell(r)=r+r^{-1}$.
Theorem 5.3.
The spectral decomposition of $(\bm{L},\tilde{\bm{L}})$ gives the following direct integral representation
$\mathcal{H}\simeq\sum_{m_{r}\in\mathbb{Z}/N\mathbb{Z}}\int_{0}^{\infty}\mathrm{d}\mu_{r}\,\varrho(\mu_{r},m_{r})^{-1}\,\mathcal{H}_{\mu_{r},m_{r}}.$ (5.32)